jhanell123:

Tom and Jack spend a certain amount of money from their accounts each week at a pet shelter. The table shows the relationship between the amount of money (y) remaining in Tom's account and the number of weeks (x): Function 1: Number of Weeks (x) Amount Remaining (dollars) (y) 1 60 2 53 3 46 4 39 The equation shows the relationship between the amount of money, y, remaining in Jack's account and the number of weeks, x: Function 2: y = −8x + 60 Which statements states and explains which function shows a greater rate of change? (1 point) a Function 1, because Tom spends $7 each week and Jack spends −$8 each week b Function 1, because Tom spends $21 each week and Jack spends $52 each week c Function 2, because Jack spends $60 each week and Tom spends $21 each week d Function 2, because Jack spends $8 each week and Tom spends $7 each week

PDOfficerLaw:

@pdofficerlaw wrote:
Each of them starts at 60. \[y^1=53 ~~~~~~~~~~x^1=2\] \[y^2=60 ~~~~~~~~~~x^2=1\] \[ROC=SLOPE=(y^2-y^1)(x^2-x^1)\] ROC=Rate Of Change
Remark: \[ROC=SLOPE=(y^2-y^1)/(x^2-x^1)\]

BrokenMoon:

no direct answers allowed

PDOfficerLaw:

@chrisdjackson wrote:
D) Function 2, because Jack spends $8 each week and Tom spends $7 each week
Are you at least going to explain it? This is considered a direct answer and is against the QC rules.

BrokenMoon:

@pdofficerlaw wrote:
@chrisdjackson wrote:
D) Function 2, because Jack spends $8 each week and Tom spends $7 each week
Are you at least going to explain it? This is considered a direct answer and is against the QC rules.
yours is consdiered a direct answer as well _-_

PDOfficerLaw:

I used the numbers from the problem and gave the user an equation to solve the question, the post is simply an equation and the numbers that need to be installed to find out the answer.